In the first year Ashok's salary was increased by 15%.In second year the salary increased by 20%.In third year the salary increased by 25%.If it is increased by 205 every year,Ashok would have got Rs.600 less than today's salary.What was Ashok's salary before first year?
Answer by stanbon(57347)

In the first year Ashok's salary was increased by 15%.In second year the salary increased by 20%.In third year the salary increased by 25%.If it is increased by 205 every year,Ashok would have got Rs.600 less than today's salary.What was Ashok's salary before first year?
--------------

Let "x" be the first year salary.
I am going to assume that "today's salary" means his salary
after all the raises.
EQUATION:
x(1.15)(1.20)(1.25) =x+3(205)+600
1.725x = x+615+600
1.725x = x+1215
0.725x = 1215
x=$1,675.86
Cheers,
Stan H.